# Probing Self-Interacting Dark Matter via Gravitational-Wave Background from Eccentric Supermassive Black Hole Mergers

### v1

@@ IV CHARACTERISTIC STRAIN OF GRAVITATIONAL WAVES
The details of the polynomials are elaborated in the Appendix. C.
For eccentric orbits, the frequency of GWs is no longer simply twice the orbital frequency. For each source, we calculate the contributions from all orbital harmonics [106]
@@
The value of the semi-major axis $a^{3}=GM_{\mathrm{tot}}/(2\pi f^{n}_{\mathrm{orb}})$ is given by Kepler’s law, and $g(n,e)$ is defined as [106]
@@
In the right panel we consider a uniformly distributed initial eccentricity $e_{0}=U\left[0.01,0.99\right]$, and we average over all eccentricities to obtain the characteristic strain curves of the SGWB corresponding to different parameters in the right figure. Note that the gray dashed line represents the characteristic strain spectrum for purely circular orbits without environmental effects. It can be observed that due to higher transition velocities $v_{t}$ and self-interaction $\sigma_{0}/m_{\chi}$ leading to lower spike densities, the dynamical friction effect is weakened, resulting in higher characteristic strain.
@@ V PARAMETER ESTIMATION FOR THE DARK PHOTON MODEL
$-2\ln\mathcal{L}(\Theta)=\sum_{\begin{subarray}{c}\{\text{EPTA, PPTA}\\ \text{NANOGrav}\}\end{subarray}}\left[\frac{h_{c,i}-h_{c}(f_{i};\Theta)}{\sigma_{i}}\right]^{2}.$ (35)
@@
The selection of the parameter prior ranges is based on the self-consistency of this model as well as previous observational or simulation studies. The NANOGrav dataset demonstrates that $\psi_{0}$ values in the range of $-3$ to $-1.5$ are physically plausible when accounting for astrophysical environmental effects [51].
As mentioned in Sec. II, the transition speed $v_{t}$ is proportional to the mass ratio between mediator and dark matter. Based on $N$-body simulations of SIDM halos and observational data of dark halos in dwarf galaxies, galaxies, and galaxy clusters, it is shown that the mass ratio is typically on the order of $10^{-3}$ [19, 113]. Additionally, some studies suggest that for SIDM, assuming a constant cross section, the numerical limit typically ranges from $0.1-10\,\mathrm{cm}^{2}/\mathrm{g}$ [20, 114, 113, 115, 116, 117]. In the environment of supermassive black hole mergers considered in this paper, where a velocity-dependent cross section is considered, the cross section is generally required to be less than $1-2\,\mathrm{cm}^{2}/\mathrm{g}$ when the dark matter particle velocity reaches $1000\,\mathrm{km}/\mathrm{s}$ or even higher [118, 119, 120].
@@
The posterior results of the parameter estimation are presented in Fig. 4. The blue region shows the posterior based on current $h_{c,i}$ uncertainties, while the orange region is the projection when these uncertainties are halved in future. We integrated the contribution of initial eccentricity with a uniform distribution, averaging the parameter estimation offsets in dark matter models across different initial eccentricities. We have also computed a scenario with a high‐eccentricity normal distribution for initial eccentricity as shown in Fig. 7.
In our eccentricity-averaged posterior results, the posterior prefers the transition velocity $v_{t}$ around the $1\sigma$ interval $356.06^{+196.9}_{-131.3}$ ($68\%$ confidence level (CL)), and the normalization constant $\psi_{0}$ to be $-2.31^{+0.11}_{-0.11}$. Note that the parameter $\sigma_{0}/m_{\chi}$ does not represent the actual interaction cross section, but requires rescaling to $1\,\mathrm{Gyr}$ for comparison across different core ages, yielding an actual range of $\sigma_{0}/m_{\chi}\cdot(t_{\mathrm{age}}/1\,\mathrm{Gyr})\simeq 5.24^{+3.25}_{-3.03}$. Nonetheless, $\sigma_{0}/m_{\chi}$ is only weakly favored due to its relatively smaller impact on $h_{c}$, and as the PTAs data errors decrease, the inference improves such that a distinct peak emerges in its posterior distribution.
To ensure the reliability of our model, we also calculated the chi-square values between the theoretical predictions and observational data. Using our best-fit parameters $v_{t}=396.9\,\mathrm{km\,s^{-1}}$, $\sigma_{0}/m_{\chi}=2.82\,\mathrm{cm^{2}\,g^{-1}}$, and $\psi_{0}=-2.31$, and the lowest five frequency bins from the NANOGrav 15-year dataset, ten from PPTA, and six from EPTA, we obtain a minimum chi-square $\chi^{2}_{\mathrm{min}}=15.588$. Additionally, adopting the parameter set corresponding to the central values from the posterior distributions gives a chi-square $\chi^{2}=16.344$. Compared to the scenario without dark matter or when considering eccentric orbital evolution alone, which results in $\chi^{2}_{\mathrm{GW}}=21.58$, our best-fit and central posterior scenarios exhibit sizable improvements.
Table 1: The table summarizes the posterior estimates of the transition velocity $v_{t}$ and the cross-section normalization $\sigma_{0}/m_{\chi}$ in SIDM model, evaluated at a core age of $1\,\mathrm{Gyr}$ for different assumed initial eccentricities. Values marked with the $\diamondsuit$ symbol correspond to the case where the uncertainties in the PTA data are reduced by a factor of two in future. The last two rows present posterior estimates obtained by averaging over contributions from different initial eccentricities and by assuming a normal distribution for the initial eccentricity, respectively.
@@
Figure 5: Upper: The posterior distribution of the interaction cross-section at a relative dark matter particle velocity $v_{\mathrm{rel}}=1500\,\mathrm{km/s}$, obtained by averaging over contributions from initial eccentricities. The green dashed line represents the central value, while the blue dashed lines indicate the ±$\pm 1\sigma$ confidence level (CL) intervals. Lower: The posterior distribution of the interaction cross-section at relative velocity $v_{\mathrm{rel}}=1500\,\mathrm{km/s}$ obtained with fixed initial eccentricity $e_{0}=0.9$.
We present in Table. 1 the posterior estimates of the transition velocity and cross-section constant under different initial eccentricity conditions. For higher initial eccentricities, posterior analysis shows increased central values of both the transition velocity and interaction cross-section constant, consistent with our model’s prediction that eccentric orbits enhance gravitational radiation-induced orbital decay. The inferred mediator to dark matter mass ratio $m_{A^{\prime}}/m_{\chi}=v_{t}$ grows monotonically with eccentricity, implying a range $\left(m_{A^{\prime}}/m_{\chi}\right)\lesssim 0.224\%$ for SIDM scenarios.
Meanwhile, in Fig. 5, we present the posterior distribution of the velocity-dependent cross section obtained from the posterior distributions of $v_{t}$ and $\sigma_{0}/m_{\chi}$, evaluated at a reference dark matter particle relative velocity of $v_{\mathrm{rel}}=1500\,\mathrm{km}/\mathrm{s}$. With a uniform initial eccentricity, averaging over all orbits gives $\sigma_{0}/m_{\chi}\simeq 0.054\,\mathrm{cm^{2}\,g^{-1}}$ for a core age of $1\,\mathrm{Gyr}$. For highly eccentric binaries ($e_{0}=0.9$), the larger inferred $v_{t}$ raises the estimate to $\sigma(v)/m_{\chi}\simeq 0.664\,\mathrm{cm^{2}\,g^{-1}}$, setting an upper bound for SIDM model.
In Fig. 6, we show the estimated range of $\sigma(v)/m_{\chi}$ at different initial eccentricities and compare it with observational data on SIDM cross sections. We adopt a mean dark matter particle relative velocity of $1500\,\mathrm{km}/\mathrm{s}$ and $2000\,\mathrm{km}/\mathrm{s}$. Notably, the estimated cross section exhibits a monotonic increase with orbital eccentricity, a trend independent of whether initial eccentricity drives the orbit toward extreme values. We also predict in the lower panel the distribution obtained when the error bars are halved in future and the estimated range becomes noticeably narrower as the uncertainties decrease.
@@
Additionally, Table 2 presents a comparative analysis of SIDM cross section limits derived under different reference velocities, benchmarked against existing observational and simulation-based studies. It shows we are able to offer new ranges for the cross section within corresponding velocity intervals for both dwarf galaxies and galaxy clusters. These ranges can also indicate upper limits on the interaction cross section derived under the framework of eccentric orbital dynamics. Furthermore, these results demonstrate the feasibility of utilizing PTAs data to investigate the environment surrounding supermassive black holes. Particularly for various dark matter models, novel parameter ranges could be identified by quantifying deviations between the characteristic strain of the SGWB and model predictions.
Figure 6: The upper figure shows violin plots of the posterior distribution of cross sections for dark matter particle relative velocities $v_{\mathrm{rel}}=1500$ and $v_{\mathrm{rel}}=2000\,\mathrm{km/s}$ under different fixed initial eccentricities. The error bars represent the central values and ±$\pm 1\sigma$ CL intervals of the cross section posterior distributions. The lower figure displays violin plots of the projected cross section posterior distributions when the error of characteristic strain values from PTA data are reduced by half in future. Compared to the upper panel, the corresponding error bar intervals become narrower. Both plots clearly demonstrate the influence of initial eccentricity values on cross section estimation, and show that the estimated values increase monotonically with increasing initial eccentricity.
Table 2: Based on the fit ranges we obtained for the self-interacting cross section and transition velocity, we calculate the projected ranges the cross section at particle velocities corresponding to the following astrophysical scales and compare them with previously established bounds.
@@
Stellar dispersion, lensing [114, 19] $\sim 1\,\mathrm{cm}^{2}/\mathrm{g}$ $0.664^{+0.708}_{-0.355}\,\mathrm{cm}^{2}/\mathrm{g}$ $1500\,\mathrm{km}/\mathrm{s}$
DM-galaxy offset (Abell 3827) [113] $\sim 1.5\,\mathrm{cm}^{2}/\mathrm{g}$ $\left(\frac{t_{\mathrm{age}}}{1\,\mathrm{Gyr}}\right)\cdot 0.664^{+0.708}_{-0.355}\,\mathrm{cm}^{2}/\mathrm{g}$ $1500\,\mathrm{km}/\mathrm{s}$
DM-galaxy offset (Abell 520) [115, 116, 117] $\sim 2\,\mathrm{cm}^{2}/\mathrm{g}$ $\left(\frac{t_{\mathrm{age}}}{1\,\mathrm{Gyr}}\right)\cdot 0.086^{+0.092}_{-0.046}\,\mathrm{cm}^{2}/\mathrm{g}$ $\sim 2500\,\mathrm{km}/\mathrm{s}$
Cluster lensing surveys (Ellipticity) [118] $\leq 1\,\mathrm{cm}^{2}/\mathrm{g}$ $\left(\frac{t_{\mathrm{age}}}{1\,\mathrm{Gyr}}\right)\cdot 1.176^{+1.254}_{-0.628}\,\mathrm{cm}^{2}/\mathrm{g}$ $1300\,\mathrm{km}/\mathrm{s}$
DM-galaxy offset (Substructure mergers) [119, 120] $\lesssim 2\,\mathrm{cm}^{2}/\mathrm{g}$ $\left(\frac{t_{\mathrm{age}}}{1\,\mathrm{Gyr}}\right)\cdot 0.085^{+0.046}_{-0.029}\,\mathrm{cm}^{2}/\mathrm{g}$ $500-4000\,\mathrm{km}/\mathrm{s}$
Mass-to-light ratio (Bullet Cluster) [122] $<0.7\,\mathrm{cm}^{2}/\mathrm{g}$ $\left(\frac{t_{\mathrm{age}}}{1\,\mathrm{Gyr}}\right)\cdot 0.013^{+0.014}_{-0.007}\,\mathrm{cm}^{2}/\mathrm{g}$ $4000\,\mathrm{km}/\mathrm{s}$
@@ VI Conclusion
This work adopts an eccentric orbital supermassive black hole binary (SMBHB) merger model surrounded by self-interacting dark matter (SIDM) and incorporates additional drag effects in dynamical friction to calculate the binary’s evolution and gravitational-wave (GW) emission, thereby obtaining the corresponding stochastic gravitational wave (SGWB) spectrum. By comparing with the characteristic strain data of SGWB from NANOGrav, PPTA, and EPTA, we establish a Bayesian model and determine the current nHz-band GW detection’s preferred range for SIDM model parameters.
Considering a uniform initial eccentricity distribution, the analysis yields preferred ranges for the mediator to dark matter mass ratio as $m_{A^{\prime}}/m_{\chi}\approx 0.119^{+0.006}_{-0.004}\%$ and for the cross section as $\sigma(v)/m_{\chi}\cdot(t_{\mathrm{age}}/1\,\mathrm{Gyr})\approx 0.054^{+0.251}_{-0.044}\,\mathrm{cm}^{2}/\mathrm{g}$ at $v_{\mathrm{rel}}=1500\,\mathrm{km/s}$. Furthermore, we demonstrate that the initial eccentricity influences the cross section estimation significantly, and our results suggest a preferred range of the mediator-to-dark matter mass ratio and the cross section such that $m_{A^{\prime}}/m_{\chi}\lesssim 0.22\%$ and $\sigma(v)/m_{\chi}\cdot(t_{\mathrm{age}}/1\,\mathrm{Gyr})\lesssim 0.66\,\mathrm{cm}^{2}/\mathrm{g}$ at $v_{\mathrm{rel}}=1500\,\mathrm{km/s}$. These findings demonstrate the potential of SGWB observations to probe dark matter particle properties. We anticipate that the improvement in the precision of future PTAs data will lead to more accurate estimation of dark matter model parameters.

### v2

@@ IV CHARACTERISTIC STRAIN OF GRAVITATIONAL WAVES
The details of the polynomials are elaborated upon in Appendix C.
For eccentric orbits, the frequency of GWs is no longer simply twice the orbital frequency. For each source, we calculate the contributions from all orbital harmonics [107]:
@@
The value of the semimajor axis $a^{3}=GM_{\mathrm{tot}}/(2\pi f^{n}_{\mathrm{orb}})$ is given by Kepler’s law, and $g(n,e)$ is defined as [107]
@@
In the right panel we consider a uniformly distributed initial eccentricity $e_{0}=U\left[0.01,0.99\right]$, and we average over all eccentricities to obtain the characteristic strain curves of the SGWB corresponding to different parameters in the right figure. Except for the gray curve that represents the characteristic strain spectrum for purely circular orbits without environmental effects, we fix the normalization parameter $\psi_{0}=-2.31$ for all spectra to facilitate a direct comparison of the spectral shapes. It can be observed that due to higher transition velocities $v_{t}$ and self-interaction $\sigma_{0}/m_{\chi}$ leading to lower spike densities, the dynamical friction effect is weakened, resulting in higher characteristic strain.
@@ V PARAMETER ESTIMATION FOR THE DARK PHOTON MODEL
$2\ln\mathcal{L}(\Theta)=\sum_{\begin{subarray}{c}\{\text{EPTA, PPTA}\\ \text{NANOGrav}\}\end{subarray}}\left[\frac{h_{c,i}-h_{c}(f_{i};\Theta)}{\sigma_{i}}\right]^{2}.$ (35)
@@
The selection of the parameter prior ranges is based on the self-consistency of this model as well as previous observational or simulation studies. The NANOGrav dataset demonstrates that $\psi_{0}$ values in the range of 3-3 to 1.5-1.5 are physically plausible when accounting for astrophysical environmental effects [52].
As mentioned in Sec. II, the transition speed $v_{t}$ is proportional to the mass ratio between the mediator and dark matter. Based on $N$-body simulations of SIDM halos and observational data of dark halos in dwarf galaxies, galaxies, and galaxy clusters, it is shown that the mass ratio is typically on the order of $10^{-3}$ [19, 114]. Additionally, some studies suggest that for SIDM, assuming a constant cross section, the numerical limit typically ranges from $0.1-10\,\mathrm{cm}^{2}/\mathrm{g}$ [20, 115, 114, 116, 117, 118]. In the environment of supermassive black hole mergers considered in this paper, where a velocity-dependent cross section is considered, the cross section is generally required to be less than $1-2\,\mathrm{cm}^{2}/\mathrm{g}$ when the dark matter particle velocity reaches $1000\,\mathrm{km}/\mathrm{s}$ or even higher [119, 120, 121].
@@
The posterior results of the parameter estimation are presented in Fig. 4. The blue region shows the posterior based on current $h_{c,i}$ uncertainties, while the orange region is the projection when these uncertainties are halved in the future. We integrated the contribution of initial eccentricity with a uniform distribution, averaging the parameter estimation offsets in dark matter models across different initial eccentricities. We have also computed a scenario with a high‐eccentricity normal distribution for initial eccentricity as shown in Fig. 7.
In our eccentricity-averaged posterior results, the posterior prefers the transition velocity $v_{t}$ around the $1\sigma$ interval $356.06^{+196.9}_{-131.3}$ ($68\%$ CL), and the normalization constant $\psi_{0}$ to be $2.31^{+0.11}_{-0.11}$. Note that the parameter $\sigma_{0}/m_{\chi}$ does not represent the actual interaction cross section, but requires rescaling to $1\,\mathrm{Gyr}$ for comparison across different core ages, yielding an actual range of $\sigma_{0}/m_{\chi}\cdot(t_{\mathrm{age}}/1\,\mathrm{Gyr})\simeq 5.24^{+3.25}_{-3.03}$. Although the overall posterior distribution shows no obvious degeneracy between the cross section and the transition velocity, a trend emerges when the $\sigma_{0}/m_{\chi}$ falls below $1\,\mathrm{cm}^{2}/\mathrm{g}$: $\sigma_{0}/m_{\chi}$ and $v_{t}$ vary inversely. This occurs because, in this low-$\sigma_{0}/m_{\chi}$ regime, the impact of $\sigma_{0}/m_{\chi}$ on $h_{c}$ becomes comparable to that of the transition velocity. And as the PTA data errors decrease, the inference improves such that a distinct peak emerges in its posterior distribution.
To ensure the reliability of our model, we also calculated the chi-square values between the theoretical predictions and observational data. Using our best-fit parameters $v_{t}=396.9\,\mathrm{km\,s^{-1}}$, $\sigma_{0}/m_{\chi}=2.82\,\mathrm{cm^{2}\,g^{-1}}$, and $\psi_{0}=-2.31$, and the lowest five frequency bins from the NANOGrav 15-year dataset, ten from PPTA, and six from EPTA, we obtain a minimum chi square $\chi^{2}_{\mathrm{min}}=15.588$. Additionally, adopting the parameter set corresponding to the central values from the posterior distributions gives a chi square $\chi^{2}=16.344$. Compared to the scenario without dark matter or when considering eccentric orbital evolution alone, which results in $\chi^{2}_{\mathrm{GW}}=21.58$, our best-fit and central posterior scenarios exhibit sizable improvements.
Table 1: The table summarizes the posterior estimates of the transition velocity $v_{t}$ and the cross section normalization $\sigma_{0}/m_{\chi}$ in the SIDM model, evaluated at a core age of $1\,\mathrm{Gyr}$ for different assumed initial eccentricities. Values marked with the $\diamondsuit$ symbol correspond to the case where the uncertainties in the PTA data are reduced by a factor of 2 in the future. The last two rows present posterior estimates obtained by averaging over contributions from different initial eccentricities and by assuming a normal distribution for the initial eccentricity, respectively.
@@
Figure 5: Upper: the posterior distribution of the interaction cross section at a relative dark matter particle velocity $v_{\mathrm{rel}}=1500\,\mathrm{km/s}$, obtained by averaging over contributions from initial eccentricities. The green dashed line represents the central value, while the blue dashed lines indicate the $\pm 1\sigma$ confidence level (CL) intervals. Lower: the posterior distribution of the interaction cross section at relative velocity $v_{\mathrm{rel}}=1500\,\mathrm{km/s}$ obtained with fixed initial eccentricity $e_{0}=0.9$.
We present in Table 1 the posterior estimates of the transition velocity and cross section constant under different initial eccentricity conditions. Physically we anticipate that both parameters should increase monotonically with eccentricity. The current data, however, are insufficient to substantiate this expectation—indeed, when the error bars are reduced, the estimated cross section decreases. This discrepancy may stem from the limited precision of the present data and from the differing spectral sensitivities and partial degeneracy between $v_{t}$ and $\sigma_{0}/m_{\chi}$. The inferred mediator to the dark matter mass ratio $m_{A^{\prime}}/m_{\chi}=v_{t}$ grows monotonically with eccentricity, implying a range of $\left(m_{A^{\prime}}/m_{\chi}\right)\lesssim 0.224\%$ for SIDM scenarios.
Meanwhile, in Fig. 5, we present the posterior distribution of the velocity-dependent cross section obtained from the posterior distributions of $v_{t}$ and $\sigma_{0}/m_{\chi}$, evaluated at a reference dark matter particle relative velocity of $v_{\mathrm{rel}}=1500\,\mathrm{km}/\mathrm{s}$. With a uniform initial eccentricity, averaging over all orbits gives $\sigma_{0}/m_{\chi}\simeq 0.054\,\mathrm{cm^{2}\,g^{-1}}$ for a core age of $1\,\mathrm{Gyr}$. For highly eccentric binaries ($e_{0}=0.9$), the larger inferred $v_{t}$ raises the estimate to $\sigma(v)/m_{\chi}\simeq 0.664\,\mathrm{cm^{2}\,g^{-1}}$, setting an upper bound for the SIDM model.
In Fig. 6, we show the estimated range of $\sigma(v)/m_{\chi}$ at different initial eccentricities. We adopt a mean dark matter particle relative velocity of $1500\,\mathrm{km}/\mathrm{s}$ and $2000\,\mathrm{km}/\mathrm{s}$. Notably, the estimated cross section exhibits a monotonic increase with orbital eccentricity, a trend independent of whether initial eccentricity drives the orbit toward extreme values. We also predict in the lower panel the distribution obtained when the error bars are halved in the future and the estimated range becomes noticeably narrower as the uncertainties decrease.
@@
Additionally, Table 2 presents a comparative analysis of SIDM cross section limits derived under different reference velocities, benchmarked against existing observational and simulation-based studies. It shows we are able to offer new ranges for the cross section within corresponding velocity intervals for both dwarf galaxies and galaxy clusters. These ranges can also indicate upper limits on the interaction cross section derived under the framework of eccentric orbital dynamics. Furthermore, these results demonstrate the feasibility of utilizing PTA data to investigate the environment surrounding supermassive black holes. Particularly for various dark matter models, novel parameter ranges could be identified by quantifying deviations between the characteristic strain of the SGWB and model predictions.
Figure 6: The upper figure shows violin plots of the posterior distribution of cross sections for dark matter particle relative velocities $v_{\mathrm{rel}}=1500$ and $v_{\mathrm{rel}}=2000\,\mathrm{km/s}$ under different fixed initial eccentricities. The error bars represent the central values and $\pm 1\sigma$ CL intervals of the cross section posterior distributions. The lower figure displays violin plots of the projected cross section posterior distributions when the error of characteristic strain values from the PTA data are reduced by half in the future. Compared to the upper panel, the corresponding error bar intervals become narrower. Both plots clearly demonstrate the influence of initial eccentricity values on cross section estimation, and show that the estimated values increase monotonically with increasing initial eccentricity.
Table 2: Based on the fit ranges we obtained for the self-interacting cross section and transition velocity, we calculate the projected ranges of the cross section at particle velocities corresponding to the following astrophysical scales and compare them with previously established bounds.
@@
Stellar dispersion,
lensing [115, 19]
$\sim 1\,\mathrm{cm}^{2}/\mathrm{g}$ $0.664^{+0.708}_{-0.355}\,\mathrm{cm}^{2}/\mathrm{g}$ $1500\,\mathrm{km}/\mathrm{s}$
DM-galaxy offset
(Abell 3827) [114]
$\sim 1.5\,\mathrm{cm}^{2}/\mathrm{g}$ $\left(\frac{t_{\mathrm{age}}}{1\,\mathrm{Gyr}}\right)\cdot 0.664^{+0.708}_{-0.355}\,\mathrm{cm}^{2}/\mathrm{g}$ $1500\,\mathrm{km}/\mathrm{s}$
DM-galaxy offset
(Abell 520) [116, 117, 118]
$\sim 2\,\mathrm{cm}^{2}/\mathrm{g}$ $\left(\frac{t_{\mathrm{age}}}{1\,\mathrm{Gyr}}\right)\cdot 0.086^{+0.092}_{-0.046}\,\mathrm{cm}^{2}/\mathrm{g}$ $\sim 2500\,\mathrm{km}/\mathrm{s}$
Cluster lensing surveys
(Ellipticity) [119]
$\leq 1\,\mathrm{cm}^{2}/\mathrm{g}$ $\left(\frac{t_{\mathrm{age}}}{1\,\mathrm{Gyr}}\right)\cdot 1.176^{+1.254}_{-0.628}\,\mathrm{cm}^{2}/\mathrm{g}$ $1300\,\mathrm{km}/\mathrm{s}$
DM-galaxy offset
(Substructure mergers) [120, 121]
$\lesssim 2\,\mathrm{cm}^{2}/\mathrm{g}$ $\left(\frac{t_{\mathrm{age}}}{1\,\mathrm{Gyr}}\right)\cdot 0.085^{+0.046}_{-0.029}\,\mathrm{cm}^{2}/\mathrm{g}$ $500-4000\,\mathrm{km}/\mathrm{s}$
Mass-to-light ratio
(Bullet Cluster) [123]
$0.7\,\mathrm{cm}^{2}/\mathrm{g}$ $\left(\frac{t_{\mathrm{age}}}{1\,\mathrm{Gyr}}\right)\cdot 0.013^{+0.014}_{-0.007}\,\mathrm{cm}^{2}/\mathrm{g}$ $4000\,\mathrm{km}/\mathrm{s}$
@@ VI Conclusion
This work adopts an eccentric orbital SMBHB merger model surrounded by SIDM and incorporates additional drag effects in dynamical friction to calculate the binary’s evolution and GW emission, thereby obtaining the corresponding SGWB spectrum. By comparing with the characteristic strain data of SGWB from NANOGrav, PPTA, and EPTA, we establish a Bayesian model and determine the current nHz-band GW detection’s preferred range for SIDM model parameters.
Considering a uniform initial eccentricity distribution, the analysis yields preferred ranges for the mediator to dark matter mass ratio as $m_{A^{\prime}}/m_{\chi}\approx 0.119^{+0.006}_{-0.004}\%$ and for the cross section as $\sigma(v)/m_{\chi}\cdot(t_{\mathrm{age}}/1\,\mathrm{Gyr})\approx 0.054^{+0.251}_{-0.044}\,\mathrm{cm}^{2}/\mathrm{g}$ at $v_{\mathrm{rel}}=1500\,\mathrm{km/s}$. Furthermore, we demonstrate that the initial eccentricity influences the cross section estimation significantly, and our results suggest a preferred range of the mediator-to-dark matter mass ratio and the cross section such that $m_{A^{\prime}}/m_{\chi}\lesssim 0.22\%$ and $\sigma(v)/m_{\chi}\cdot(t_{\mathrm{age}}/1\,\mathrm{Gyr})\lesssim 0.66\,\mathrm{cm}^{2}/\mathrm{g}$ at $v_{\mathrm{rel}}=1500\,\mathrm{km/s}$. These findings demonstrate the potential of SGWB observations to probe dark matter particle properties. We anticipate that the improvement in the precision of future PTAs data will lead to a more accurate estimation of dark matter model parameters.